OpenStudy (anonymous):

If the second equation is multiplied by 4 and then the equations are added, the result is _____. x + 4y = 1 x - y = -3

OpenStudy (bahrom7893):

x+4y = 1 4x-4y = -12 ------------- 5x + 0 = -11 or 5x = -11